OpenStudy (lilai3):

Jolon will be staying with his grandparents for a week. There are 4 flights that leave the airport near Jolon's home that connect to an airport that has 2 different flights to his granparents' hometown. Find the number of possible flights. Then find the probability of taking the earliest flight from each airport if the flight is selected at random.

OpenStudy (dumbcow):

4*2 = 8 probability = 1/8

OpenStudy (mani_jha):

For the first trip, he can choose any of the four flights, and for each of these four possible flights, he can choose any two flights for the next trip. So, no of possible flights=4*2=8

OpenStudy (anonymous):

Assuming they all connect, dumbcow beat me to it. Need to work on reading and typing faster, LOL.

OpenStudy (lilai3):

~thanks, guys!~

OpenStudy (anonymous):

The probability is one eighth because there is only one of the combinations (of the eight possible ways to get there) that meet the specification of being the first from each airport.

OpenStudy (lilai3):

:D i new i could devend on vu! >:V